# Constraining interacting dark energy models with black hole superradiance

### v1

@@ Appendix A Derivation of the effective energy transfer rate $Q$ from field theory
$\dot{\rho}_{\phi}+3H\left(\rho_{\phi}+P_{\phi}\right)=-Q\,$, (A.15)
$\dot{\rho}_{\mathrm{dm}}+3H\rho_{\mathrm{dm}}=Q\,$, (A.16)
@@
$Q=\frac{\Omega^{\prime}(\Phi)}{\Omega(\Phi)}\dot{\Phi}\rho_{\mathrm{dm}}\,.$ (A.17)
@@ Appendix B Effective potential in the non-minimally coupled quintessence model
$S=\int\mathrm{d}^{4}x\sqrt{-g^{E}}\left[\frac{R^{E}}{16\pi G}-\frac{1}{2}g_{\mu\nu}^{E}\partial^{\mu}\Phi\partial^{\nu}\Phi-{V(\Phi)}\right]+S\left(\Psi;\Omega^{2}(\Phi)g_{\mu\nu}^{E}\right)\,.$ (B.1)
@@
$G_{\mu\nu}^{E}$ $=8\pi G\left(T_{\mu\nu}^{E}+\partial_{\mu}\Phi\partial_{\nu}\Phi-\frac{g_{\mu\nu}^{E}}{2}(\partial\Phi)^{2}-{g_{\mu\nu}^{E}}V(\Phi)\right)$ (B.2)
@@
${T}^{J}_{\mu\nu}=\frac{-2}{\sqrt{-{g}^{J}}}\frac{\partial}{\partial{g}^{J,\mu\nu}}\left(\sqrt{-{g}^{J}}\,\mathcal{L}_{\mathrm{dm}}\left[{g}^{J}_{\mu\nu},\Psi\right]\right)\,.$ (B.4)
@@
$\nabla^{E}_{\mu}T_{\nu}^{E,\mu}=T^{E}\Omega^{-1}\nabla_{\nu}\Omega\,.$ (B.5)
@@
$\nabla_{t}\left(\Omega^{3w-1}\rho\right)=0\,.$ (B.6)
@@
$\square^{E}\Phi=(1-3w)\Omega^{\prime}\Omega^{-3w}\hat{\rho}+{V^{\prime}(\Phi)}\,.$ (B.9)
@@
$V_{\mathrm{eff}}(\Phi)=V(\Phi)+\Omega^{1-3w}(\Phi)\hat{\rho}\,$. (B.10)
For pressureless dark matter ($w=0$), this simplifies to $V_{\mathrm{eff}}(\Phi)=V(\Phi)+\Omega(\Phi)\hat{\rho}_{\mathrm{dm}}$ as shown in Section 5.
@@ Appendix C Relating DM halo parameters to supermassive-black-hole mass
When a BH surrounded by $\rho(r)=\rho_{0}(r/r_{0})^{-\gamma}$ DM profile grows adiabatically, it gravitationally concentrates the surrounding DM around it to a cuspy profile usually referred to as a “DM spike”. The DM density spike can be modelled generally as [PhysRevLett.83.1719]
$\rho_{\mathrm{sp}}(r)=\rho_{R}\Theta(r-4R_{s})\left(1-\frac{4R_{\mathrm{s}}}{r}\right)^{3}\left(\frac{R_{\mathrm{sp}}}{r}\right)^{\gamma_{\mathrm{sp}}}.$ (C.1)
This Appendix details the procedure, following [Nishikawa:2017chy], used to estimate the DM profile halo parameter $\rho_{0}$ and $r_{0}$ associated with a given SMBH mass $M$. These halo parameters are necessary to calculate the structure of the DM spike, specifically $\rho_{R}$ and $R_{\mathrm{sp}}$, as described in Section 5. The method assumes the DM halo follows an NFW profile outside the region influenced by the SMBH, extending to the virial radius $r_{\mathrm{vir}}$.
First, we utilize the empirical $M$-$\sigma$ relationship, which connects the SMBH mass $M$ to the one-dimensional velocity dispersion $\sigma$ of the host bulge or halo [Gebhardt:2000fk, Ferrarese:2000se] via
$\log_{10}\left(M/\mathrm{M}_{\odot}\right)=a+b\log_{10}\left(\sigma/200\mathrm{~km}\mathrm{~s}^{-1}\right)$ (C.2)
with empirically determined parameters $a=8.12$ and $b=4.24$. Second, for an NFW halo, the velocity dispersion $\sigma$ can be related to the halo parameters $\rho_{0}$ and $r_{0}$ via [Nishikawa:2017chy]
$\sigma^{2}=\frac{4\pi G_{N}\rho_{0}r_{0}^{2}g\left(c_{m}\right)}{c_{m}},$ (C.3)
@@
$M_{\mathrm{vir}}=\rho_{\mathrm{crit}}\Delta\left(\frac{4\pi\left(c\left(M_{\mathrm{vir}}\right)r_{0}\right)^{3}}{3}\right)=4\pi\rho_{0}r_{0}^{3}g\left(c\left(M_{\mathrm{vir}}\right)\right),$ (C.4)
@@
Fourth, the concentration parameter $c$ is known to correlate with the virial mass $M_{\mathrm{vir}}$ through empirically calibrated relations, often expressed as a polynomial in $\log(M_{\mathrm{vir}})$ as [Nishikawa:2017chy]
$c\left(M_{\mathrm{vir}}\right)$ $=\sum_{i=0}^{5}c_{i}\left[\log_{10}\left(\frac{hM_{\mathrm{vir}}}{\mathrm{M}_{\odot}}\right)\right]^{i},$ (C.5)
where $h$ is the dimensionless Hubble parameter, $h_{0}=0.67$ and the coefficients $c_{i}$ are fitted from simulations, $c_{0}=37.5153,c_{1}=-1.5093,c_{2}=1.636\times 10^{-2},c_{3}=3.66\times 10^{-4},c_{4}=-2.89237\times 10^{-5}$ and $c_{5}=5.32\times 10^{-7}$. Using the above equations, we obtain a relation to connect $M$ and $M_{\mathrm{vir}}$ via
$\sigma^{2}\left(M\right)=\frac{4\pi}{3}G_{N}\Delta\rho_{c}\frac{g\left(c_{\mathrm{m}}\right)c}{g(c)c_{\mathrm{m}}}\left(\frac{3M_{\mathrm{vir}}}{4\pi\Delta\rho_{c}}\right)^{2/3}.$ (C.6)
We can then solve for $M_{\mathrm{vir}}$ numerically for each SMBH mass $M$ using (C.6), and subsequently obtain $\rho_{0}$ and $r_{0}$ using (C.4). With the halo parameters $\rho_{0}$ and $r_{0}$ thus determined for a given SMBH mass $M$, the spike parameters $\rho_{R}$ and $R_{\mathrm{sp}}$ can be calculated using the formulae provided in the Section 5, allowing for the evaluation of the spike density profile $\rho_{\mathrm{sp}}(r)$.
@@ Appendix D Numerical details for solving the radial superradiance equation
This Appendix details the numerical methods used to solve the Klein-Gordon equation with a position-dependent effective mass, $\mu_{\mathrm{eff}}^{2}(r)$, as required for Model II (Section 5). The standard procedure involves separating the scalar field $\Phi(t,r,\theta,\phi)$ using the ansatz $e^{-i\omega t}e^{im\phi}R(r)S(\theta)$, leading to distinct angular and radial ordinary differential equations.
@@
$\left(1-x^{2}\right)\frac{\mathrm{d}^{2}S(x)}{\mathrm{d}x^{2}}-2x\frac{\mathrm{d}S(x)}{\mathrm{d}x}+\left(\Lambda_{\ell m}+a^{2}\omega^{2}x^{2}-\frac{m^{2}}{1-x^{2}}\right)S(x)=0\,.$ (D.1)
@@
$\lim_{x\rightarrow-1}S(x)\sim(1+x)^{|m|/2},\quad\lim_{x\rightarrow+1}S(x)\sim(1-x)^{|m|/2}\,.$ (D.2)
This equation can be solved numerically using Leaver’s continued fraction method [Leaver:1985ax, Dolan:2007mj, Liu:2022ygf]. By expanding the solution as
$S(x)=\exp(a\omega x)(1-x)^{|m|/2}(1+x)^{|m|/2}\sum_{n=0}^{\infty}b_{n}(1+x)^{n},$ (D.3)
@@
$\left\{\begin{array}[]{l}\alpha_{0}b_{1}+\beta_{0}b_{0}=0,\\ \alpha_{n}b_{n+1}+\beta_{n}b_{n}+\gamma_{n}b_{n-1}=0\,,\quad(n\geq 1)\end{array}\right$. (D.4)
@@
$\left\{\begin{aligned} \alpha_{n}&=-2(n+1)(n+|m|+1)\,,\\ \beta_{n}&=n(n+1)+a^{2}\omega^{2}-2a\omega m+|m|(|m|+1)-\Lambda_{\ell m}+2a\omega(n+|m|)\\ &\quad-(2n+1)(a\omega+|m|)\,,\\ \gamma_{n}&=2a\omega(n+|m|)\,.\end{aligned}\right$. (D.5)
@@
$0=\beta_{0}-\displaystyle\frac{\alpha_{0}\gamma_{1}}{\beta_{1}-\displaystyle\frac{\alpha_{1}\gamma_{2}}{\beta_{2}-\displaystyle\frac{\alpha_{2}\gamma_{3}}{\beta_{3}-\ldots}}}\,.$ (D.6)
@@
However, the radial equation Eq. (5.9) is modified by the position-dependent effective mass $\mu_{\mathrm{eff}}^{2}(r)$ and cannot be solved using the same continued fraction technique. We therefore employ a direct integration method, specifically the “shooting method” [Molina:2010fb] to find the complex eigenfrequency $\omega$.
@@
$R^{\mathrm{horizon}}(r)$ $=(r-r_{+})^{\sigma_{+}}\sum_{i=0}^{\infty}h_{\mathrm{H}}^{(i)}(r-r_{+})^{i}\,$, (D.7)
$R^{\infty}(r)$ $=e^{ik_{\infty}r}r^{\gamma-1}\sum_{i=0}^{\infty}h_{\infty}^{(i)}r^{-i}\,$, (D.8)
@@
Care must be taken when implementing the numerical integration from spatial infinity. The general solution at large $r$ contains both exponentially decaying (physical) and exponentially growing (unphysical) modes. To find the correct quasi-bound state, we must select the purely decaying solution. Numerical errors can inadvertently introduce contributions from the growing mode, which can dominate the solution if the integration starts from too large a radius (“numerical infinity”). This issue can be mitigated by choosing a sufficiently small value for numerical infinity (e.g., $r_{\infty}\sim 50M$) and using a sufficiently high order in the asymptotic series expansion to set the initial conditions accurately [Molina:2010fb].

### v2

@@ Appendix A Derivation of the effective energy transfer rate $Q$ from field theory
$\dot{\rho}_{\phi}+3H\left(\rho_{\phi}+P_{\phi}\right)=-Q\,$, (54)
$\dot{\rho}_{\mathrm{dm}}+3H\rho_{\mathrm{dm}}=Q\,$, (55)
@@
$Q=\frac{\Omega^{\prime}(\Phi)}{\Omega(\Phi)}\dot{\Phi}\rho_{\mathrm{dm}}\,.$ (56)
@@ Appendix B Effective potential in the non-minimally coupled quintessence model
$S$ $=\int\mathrm{d}^{4}x\sqrt{-g^{E}}\left[\frac{R^{E}}{16\pi G}-\frac{1}{2}g_{\mu\nu}^{E}\partial^{\mu}\Phi\partial^{\nu}\Phi-{V(\Phi)}\right]$
$+S\left(\Psi;\Omega^{2}(\Phi)g_{\mu\nu}^{E}\right)\,.$ (57)
@@
$G_{\mu\nu}^{E}$ $=8\pi G\left(T_{\mu\nu}^{E}+\partial_{\mu}\Phi\partial_{\nu}\Phi-\frac{g_{\mu\nu}^{E}}{2}(\partial\Phi)^{2}-{g_{\mu\nu}^{E}}V(\Phi)\right)$ (58)
@@
${T}^{J}_{\mu\nu}=\frac{-2}{\sqrt{-{g}^{J}}}\frac{\partial}{\partial{g}^{J,\mu\nu}}\left(\sqrt{-{g}^{J}}\,\mathcal{L}_{\mathrm{dm}}\left[{g}^{J}_{\mu\nu},\Psi\right]\right)\,.$ (60)
@@
$\nabla^{E}_{\mu}T_{\nu}^{E,\mu}=T^{E}\Omega^{-1}\nabla_{\nu}\Omega\,.$ (61)
@@
$\nabla_{t}\left(\Omega^{3w-1}\rho\right)=0\,.$ (62)
@@
$\square^{E}\Phi=(1-3w)\Omega^{\prime}\Omega^{-3w}\hat{\rho}+{V^{\prime}(\Phi)}\,.$ (65)
@@
$V_{\mathrm{eff}}(\Phi)=V(\Phi)+\Omega^{1-3w}(\Phi)\hat{\rho}\,$. (66)
For pressureless dark matter ($w=0$), this simplifies to $V_{\mathrm{eff}}(\Phi)=V(\Phi)+\Omega(\Phi)\hat{\rho}_{\mathrm{dm}}$ as shown in Section V.
@@ Appendix C Relating DM halo parameters to supermassive-black-hole mass
When a BH surrounded by $\rho(r)=\rho_{0}(r/r_{0})^{-\gamma}$ DM profile grows adiabatically, it gravitationally concentrates the surrounding DM around it to a cuspy profile usually referred to as a “DM spike”. The DM density spike can be modelled generally as [61]
$\rho_{\mathrm{sp}}(r)=\rho_{R}\Theta(r-4R_{s})\left(1-\frac{4R_{\mathrm{s}}}{r}\right)^{3}\left(\frac{R_{\mathrm{sp}}}{r}\right)^{\gamma_{\mathrm{sp}}}.$ (67)
This Appendix details the procedure, following [94], used to estimate the DM profile halo parameter $\rho_{0}$ and $r_{0}$ associated with a given SMBH mass $M$. These halo parameters are necessary to calculate the structure of the DM spike, specifically $\rho_{R}$ and $R_{\mathrm{sp}}$, as described in Section V. The method assumes the DM halo follows an NFW profile outside the region influenced by the SMBH, extending to the virial radius $r_{\mathrm{vir}}$.
First, we utilize the empirical $M$-$\sigma$ relationship, which connects the SMBH mass $M$ to the one-dimensional velocity dispersion $\sigma$ of the host bulge or halo [58, 55] via
$\log_{10}\left(M/\mathrm{M}_{\odot}\right)=a+b\log_{10}\left(\sigma/200\mathrm{~km}\mathrm{~s}^{-1}\right)$ (68)
with empirically determined parameters $a=8.12$ and $b=4.24$. Second, for an NFW halo, the velocity dispersion $\sigma$ can be related to the halo parameters $\rho_{0}$ and $r_{0}$ via [94]
$\sigma^{2}=\frac{4\pi G_{N}\rho_{0}r_{0}^{2}g\left(c_{m}\right)}{c_{m}},$ (69)
@@
$M_{\mathrm{vir}}=\rho_{\mathrm{crit}}\Delta\left(\frac{4\pi\left(c\left(M_{\mathrm{vir}}\right)r_{0}\right)^{3}}{3}\right)=4\pi\rho_{0}r_{0}^{3}g\left(c\left(M_{\mathrm{vir}}\right)\right),$ (70)
@@
Fourth, the concentration parameter $c$ is known to correlate with the virial mass $M_{\mathrm{vir}}$ through empirically calibrated relations, often expressed as a polynomial in $\log(M_{\mathrm{vir}})$ as [94]
$c\left(M_{\mathrm{vir}}\right)$ $=\sum_{i=0}^{5}c_{i}\left[\log_{10}\left(\frac{hM_{\mathrm{vir}}}{\mathrm{M}_{\odot}}\right)\right]^{i},$ (71)
where $h$ is the dimensionless Hubble parameter, $h_{0}=0.67$ and the coefficients $c_{i}$ are fitted from simulations, $c_{0}=37.5153,c_{1}=-1.5093,c_{2}=1.636\times 10^{-2},c_{3}=3.66\times 10^{-4},c_{4}=-2.89237\times 10^{-5}$ and $c_{5}=5.32\times 10^{-7}$.
Using the above equations, we obtain a relation to connect $M$ and $M_{\mathrm{vir}}$ via
$\sigma^{2}\left(M\right)=\frac{4\pi}{3}G_{N}\Delta\rho_{c}\frac{g\left(c_{\mathrm{m}}\right)c}{g(c)c_{\mathrm{m}}}\left(\frac{3M_{\mathrm{vir}}}{4\pi\Delta\rho_{c}}\right)^{2/3}.$ (72)
We can then solve for $M_{\mathrm{vir}}$ numerically for each SMBH mass $M$ using (72), and subsequently obtain $\rho_{0}$ and $r_{0}$ using (70). With the halo parameters $\rho_{0}$ and $r_{0}$ thus determined for a given SMBH mass $M$, the spike parameters $\rho_{R}$ and $R_{\mathrm{sp}}$ can be calculated using the formulae provided in the Section V, allowing for the evaluation of the spike density profile $\rho_{\mathrm{sp}}(r)$.
@@ Appendix D Numerical details for solving the radial superradiance equation
This Appendix details the numerical methods used to solve the Klein-Gordon equation with a position-dependent effective mass, $\mu_{\mathrm{eff}}^{2}(r)$, as required for Model II (Section V). The standard procedure involves separating the scalar field $\Phi(t,r,\theta,\phi)$ using the ansatz $e^{-i\omega t}e^{im\phi}R(r)S(\theta)$, leading to distinct angular and radial ordinary differential equations.
@@
$\left(1-x^{2}\right)\frac{\mathrm{d}^{2}S(x)}{\mathrm{d}x^{2}}-2x\frac{\mathrm{d}S(x)}{\mathrm{d}x}$
$+\left(\Lambda_{\ell m}+a^{2}\omega^{2}x^{2}-\frac{m^{2}}{1-x^{2}}\right)S(x)=0\,.$ (73)
@@
$\lim_{x\rightarrow-1}S(x)\sim(1+x)^{|m|/2},\quad\lim_{x\rightarrow+1}S(x)\sim(1-x)^{|m|/2}\,.$ (74)
This equation can be solved numerically using Leaver’s continued fraction method [77, 50, 84]. By expanding the solution as
$S(x)=\exp(a\omega x)(1-x)^{|m|/2}(1+x)^{|m|/2}\sum_{n=0}^{\infty}b_{n}(1+x)^{n},$ (75)
@@
$\left\{\begin{array}[]{l}\alpha_{0}b_{1}+\beta_{0}b_{0}=0,\\ \alpha_{n}b_{n+1}+\beta_{n}b_{n}+\gamma_{n}b_{n-1}=0\,,\quad(n\geq 1)\end{array}\right$. (76)
@@
$\left\{\begin{aligned} \alpha_{n}&=-2(n+1)(n+|m|+1)\,,\\ \beta_{n}&=n(n+1)+a^{2}\omega^{2}-2a\omega m+|m|(|m|+1)-\Lambda_{\ell m}\\ &\quad+2a\omega(n+|m|)-(2n+1)(a\omega+|m|)\,,\\ \gamma_{n}&=2a\omega(n+|m|)\,.\end{aligned}\right$. (77)
@@
$0=\beta_{0}-\displaystyle\frac{\alpha_{0}\gamma_{1}}{\beta_{1}-\displaystyle\frac{\alpha_{1}\gamma_{2}}{\beta_{2}-\displaystyle\frac{\alpha_{2}\gamma_{3}}{\beta_{3}-\ldots}}}\,.$ (78)
@@
However, the radial equation Eq. (41) is modified by the position-dependent effective mass $\mu_{\mathrm{eff}}^{2}(r)$ and cannot be solved using the same continued fraction technique. We therefore employ a direct integration method, specifically the “shooting method” [91] to find the complex eigenfrequency $\omega$.
@@
$R^{\mathrm{horizon}}(r)$ $=(r-r_{+})^{\sigma_{+}}\sum_{i=0}^{\infty}h_{\mathrm{H}}^{(i)}(r-r_{+})^{i}\,$, (79)
$R^{\infty}(r)$ $=e^{ik_{\infty}r}r^{\gamma-1}\sum_{i=0}^{\infty}h_{\infty}^{(i)}r^{-i}\,$, (80)
@@
Care must be taken when implementing the numerical integration from spatial infinity. The general solution at large $r$ contains both exponentially decaying (physical) and exponentially growing (unphysical) modes. To find the correct quasi-bound state, we must select the purely decaying solution. Numerical errors can inadvertently introduce contributions from the growing mode, which can dominate the solution if the integration starts from too large a radius (“numerical infinity”). This issue can be mitigated by choosing a sufficiently small value for numerical infinity (e.g., $r_{\infty}\sim 50M$) and using a sufficiently high order in the asymptotic series expansion to set the initial conditions accurately [91].